OpenStudy (automaticloveletter):

The distances (y), in miles, of two cars from their starting points at certain times (x), in hours, are shown by the equations below: Car A y = 60x + 10 Car B y = 40x + 70 After how many hours will the two cars be at the same distance from their starting point and what will that distance be? 2 hours, 150 miles 2 hours, 190 miles 3 hours, 150 miles 3 hours, 190 miles

OpenStudy (anonymous):

hey do u still need hlp??

OpenStudy (automaticloveletter):

yes

OpenStudy (anonymous):

ok

OpenStudy (automaticloveletter):

Are you going to help?

OpenStudy (anonymous):

yess jus give me like 7minutes to work out tha problems

OpenStudy (anonymous):

What one do you think is correct?

OpenStudy (anonymous):

When you worked out the problem, what one did you come up with as being correct?

OpenStudy (anonymous):

Do you just want the answer?

OpenStudy (automaticloveletter):

Nah

OpenStudy (anonymous):

What one do you think is correct?

OpenStudy (anonymous):

I thought you did not just want the answer?

OpenStudy (automaticloveletter):

I never took it

OpenStudy (anonymous):

oh my bad i can delete and tell u how to solve it

OpenStudy (automaticloveletter):

how did you?

OpenStudy (anonymous):

well u jus enter 190 where the y is ans 3 where the x is

OpenStudy (anonymous):

\[190=60(3)+10 and 190=40(3)+70\]

OpenStudy (anonymous):

and then u solve them

OpenStudy (anonymous):

Djthereal, that works for this but she might need to know how to do it when where there are not any choices. Have fun

OpenStudy (anonymous):

lol is wat is that supposed to mean??